OpenStudy (anonymous):

solve for k. 1/8k + -6/5=4- -1/7k

OpenStudy (whpalmer4):

Can you write that with parentheses or the equation editor? It isn't clear what you mean.

OpenStudy (anonymous):

\[\frac{ 1 }{ 8 }k + \frac{ -6 }{5} = 4- \frac{ -1 }{7} \]

OpenStudy (whpalmer4):

Okay, you can try to make a common denominator, or you could do the brute force approach which is to simply multiply each term by the product of all of the denominators (8*5*7). After you do that, you won't have any fractions left, and it should be easy to find the value of k. Try it, and I'll check your answer.

OpenStudy (whpalmer4):

Oops, you dropped the k on the right hand side. \[\frac{1}{8}k + \frac{-6}{5} = 4 - \frac{-1}{7}k\]

OpenStudy (anonymous):

can you help by showing me step by step?

OpenStudy (whpalmer4):

Okay, let's multiply everything by 8*5*7: \[\frac{1}{8}k(8*5*7) + \frac{-6}{5}(8*5*7) = 4(8*5*7) - \frac{-1}{7}k(8*5*7)\]Then we cancel things out and get\[k*5*7-6*8*7 = 4*8*5*7+8*5*k\]or\[35k - 336 = 1120+ 40k\]Now subtract 35k from each side\[35k-35k-336 = 1120+40k-35k\]\[-336=1120+5k\]Subtract 1120 from each side\[-336-1120 = 1120-1120+5k\]\[5k=-1456\]Divide both sides by 5\[k=-\frac{1456}{5}\] Which is pretty strange, and makes me a bit suspicious that the problem wasn't exactly the one to be solved.

OpenStudy (whpalmer4):

But it *is* the answer to the problem we solved :-)

OpenStudy (anonymous):

thank you ;'D ure awesome.

OpenStudy (anonymous):

what did i do wrong for this next EQUATION? \[\frac{ 1 }{ 2}x + \frac{ 1 }{ 4 }= -1 (\frac{ 5 }{ 6 }x +4)\] answer i got. x =12/7

OpenStudy (whpalmer4):

And you think the correct answer is?

OpenStudy (anonymous):

x= 12/7

OpenStudy (whpalmer4):

Wait, you said you got x= 12/7, but that it was wrong.

OpenStudy (whpalmer4):

I don't think 12/7 is correct. How did you do the problem? What was your first step?

OpenStudy (anonymous):

multipyling the common denomionator 2*4*6 like the problem before.

OpenStudy (whpalmer4):

Okay, that's a good start. When I do that, I get \[24x + 12 = -40x -192\]\[64x = -204\]

OpenStudy (anonymous):

oh!!

OpenStudy (anonymous):

so then the answer would be -3.187

OpenStudy (whpalmer4):

Close, but not exactly. -3.1875.

OpenStudy (whpalmer4):

Unless a numeric answer was requested (or this is a chemistry or physics problem), I personally would leave it as a simplified fraction, -51/16.

OpenStudy (whpalmer4):

When in Rome, do as the Romans do...

OpenStudy (anonymous):

\[1(x-2)+7=-7(x+2) solve x.\]

OpenStudy (whpalmer4):

Okay, I would use the distributive property of multiplication to expand that. \[a(b+c) = ab + ac\] Then add and subtract terms from both sides to get x on one side, and a number on the other.

OpenStudy (anonymous):

x= -2/15?

OpenStudy (whpalmer4):

Well, if you put that back in the equation, do you get a true statement, or do you end up with something crazy, like 7 = 9?

OpenStudy (anonymous):

actually i gott -19/8

OpenStudy (whpalmer4):

That's better :-)

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 }x + \frac{ 1 }{3}=1(\frac{ 2 }{ 3 }x +2)\] my answer come out to be 35/3?

OpenStudy (whpalmer4):

Don't think so.

OpenStudy (anonymous):

but its 5*3*3

OpenStudy (whpalmer4):

Multiply both sides by 15, I get \[3x + 5 = 10x + 30\] which means x < 0

OpenStudy (anonymous):

-25=7x -25/7?

OpenStudy (whpalmer4):

yes, that's correct